# Graded Projection Recursion (GPR): A Framework for Controlling Bit-Complexity of Algebraic Packing

### v2

@@ 9.1 Arithmetic recurrence
Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two, then:
@@
We now consider a single call of KernelMM on size $n$.
@@ 9.2 Bit/word complexity (fixed words, staged precision)
Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers. Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
@@ 9.3 Example: integer polynomial convolution as a GPR-admissible packing
There exists a sufficient base sequence $\beta(\sigma)$ (growing with recursion depth $\sigma$) such that the recursive packed convolution algorithm recovers $f\ast g$ exactly in dyadic lattice semantics. The arithmetic recurrence is:
@@
At each node, the packed product $F\ast G$ admits the three-band form (Proposition 9.8). Dyadic normalization (Lemma 9.7) ensures the band coefficients are bounded in the normalized lattice $\mathcal{L}_{\mathrm{coeff}}(\sigma)$. By choosing $\beta$ to satisfy the sufficient-base condition (Theorem 7.1) for the current depth $\sigma$, the master gap inequality holds, and the scaled sieve $\mathcal{S}_{\beta,\sigma}$ recovers the middle band $T$ exactly. The algorithm performs $4$ subcalls (implicitly or explicitly to recover all components of the convolution from the packed forms), maintaining exactness throughout via the staged precision model.
@@ 10 Prior related work and what is new here
What is new in this note is not the existence of coefficient/band extraction, but the way recursive invariant amplification (centering plus dyadic $\ell_{2}$ normalization at every recursion node) makes coefficient extraction compatible with a model-honest intermediate-budget analysis: it gives a constant base $\beta_{\star}$, eliminates bit-growth of mantissas, and improves feasibility/robustness by bounding intermediates at every depth rather than only at the root.
@@ 10.1 APA and border-rank methods in fast matrix multiplication
In fast matrix multiplication theory, approximate (APA) bilinear algorithms introduce a formal parameter (often written $\lambda$ or $\varepsilon$) to separate contributions across different powers of that parameter and then recover a desired coefficient via an exactification step (interpolation/limit argument). This is similar to the work of Bini et al. on approximate MM [18] and in Bini’s later discussion relating exact and approximate bilinear algorithms [19]. A closely related tensor-language view is that such constructions upper-bound the border rank of the MM tensor, which is sufficient to derive asymptotic exponents (see [20] and Pan’s trilinear/field-extension methods [21]).
@@ 10.4 What changes the game: recursive invariant amplification
Part II Float-GPR: baseline comparators and commensurate accuracy (Part II)
@@ 14 Float-GPR complexity
Part III Case studies and templates (Part III)
@@ 17.2 APSP via min-plus products (semiring mode)
17.3 Triangle detection via Boolean MM (semiring mode)
Theorem 17.4 (Triangle detection via Boolean GPR kernel (sketch)).
Let $G$ be a directed or undirected graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$. Work over the Boolean semiring (,)=(,)($\oplus$$\otimes$)=($\lor$$\land$), so Boolean matrix multiplication is:
$(X\otimes Y)_{ij}~=~\bigvee_{k=1}^{n}\left(X_{ik}\wedge Y_{kj}\right).$ (171)
Assume Boolean matrix multiplication is implemented by a recursive blocked GPR kernel KernelMatMulBool that uses graded projection corresponding to word-packing of bits (i.e., each row/column block is stored as bitsets packed into machine words and block products are computed by word-level operations), then:
  1. 1.
    (Correctness) $G$ contains a triangle if and only if there exist $i,j$ such that:
    $(A\otimes A)_{ij}\wedge A_{ji}~=~1.$ (172)
    Equivalently, in Boolean arithmetic, $\mathrm{trace}(A\otimes A\otimes A)=1$ if and only if a triangle exists.
  2. 2.
    (Complexity reduction to Boolean MM) Triangle detection reduces to one Boolean matrix multiplication plus $O(n^{2})$ additional work for the final check, so the runtime is $\widetilde{O}(T_{\mathrm{BoolMM}}(n))$, where $T_{\mathrm{BoolMM}}(n)$ is the cost of one Boolean matrix multiplication implemented by KernelMatMulBool. In particular, with word-packing and a cache-friendly blocked implementation one obtains $O(n^{3}/w)$ word operations (where $w$ is the machine word size).
Proof.
Correctness
A triangle exists if and only if there are vertices $i,j,k$ with edges $(i,k)$, $(k,j)$, and $(j,i)$. The Boolean product $(A\otimes A)_{ij}=1$ holds exactly when there exists $k$ such that $A_{ik}=1$ and $A_{kj}=1$, i.e., a length-2 path from $i$ to $j$. Thus $(A\otimes A)_{ij}\wedge A_{ji}=1$ holds exactly when there is a length-2 path $i\to j$ and an edge $j\to i$, which is precisely a directed triangle. The trace characterization is equivalent.
Complexity
Computing $B:=A\otimes A$ is one Boolean matrix multiplication, and checking whether any entry satisfies $B_{ij}\wedge A_{ji}=1$ is $O(n^{2})$, so the total runtime is $\widetilde{O}(T_{\mathrm{BoolMM}}(n))$. With word-packing, each inner product reduces to a wordwise AND and an “any-nonzero” test (or popcount), giving the standard $O(n^{3}/w)$ bound. ∎
@@ 17.4 How to create new GPR reductions
Part IV Generality of GPR: admissible packings, design patterns, and extensions (Part IV)
@@ 21.4 B) Beyond rounding gaps: an extractor abstraction
Proof sketch.
Uniform master bounds imply a uniform sufficient condition for extraction once global parameters are chosen (base separation, CRT interval, bucket width, etc.). The extraction is then correct at every leaf, and correctness propagates up the recursion by the same structural induction as in Parts I–III. Complexity is unchanged asymptotically because extractor operations are $O(1)$ per entry per level (or bounded by a fixed number of moduli/buckets). ∎
@@ B.4 Exact sieve extractor $\mathcal{S}_{\beta}(\cdot)$
Fix an integer $b\geq 1$ and let $\beta:=2^{b}$. For $z\in\mathcal{L}_{D}$, define:
$\mathcal{S}_{\beta}(z)~:=~\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)-\beta\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z/\beta).$ (245)
Here $z/\beta$ is implemented by exponent update $e\leftarrow e-b$ (exact in $\mathcal{L}_{D}$), and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is implemented by Definition B.4.
Lemma B.7 (Implementation matches Part I sieve).
If $\mathcal{S}_{\beta}$ is implemented as in Definition B.6, then it matches the Part I definition $\mathcal{S}_{\beta}(z)=\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)-\beta\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z/\beta)$ exactly, and Lemmas 5.2 and 5.3 apply verbatim.
@@
All steps are exact in $\mathcal{L}_{D}$: dividing by $\beta=2^{b}$ is an exponent-offset update, and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma. Therefore the implemented $\mathcal{S}_{\beta}$ equals the mathematical sieve definition. ∎
B.5 Gap checking and failure policy (optional but recommended)
For debugging and for reference implementations, it is useful to explicitly check the master gap inequality before applying the sieve. For a packed scalar $z=\beta U+T+L/\beta$, compute the three quantities appearing in Lemma 5.3 in the lifted units and assert:
If the check fails, the implementation should either:
  1. 1.
    Return a failure flag (instance leaves the declared success set), or
  2. 2.
    Increase $\beta$ and retry (if using an adaptive base policy outside the constant-$\beta$ regime).
On the success set, the check passes and extraction is exact.

### v3

@@ 9 Complexity and implementation model
Remark 9.1 (Certified regime and staged retry).
In exact dyadic lattice semantics, Appendix  B.5 proves that the certified base-selection rule (Eq.  271 with the one-step upgrade Eq.  272) succeeds after at most one upgrade at every node, provided coefficient-lattice alignment is preserved (i.e., no mantissa bits are dropped below the local $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ resolution). Therefore, within the declared exact-execution regime, the ClassicalMM branch in Algorithm  1 is unreachable and serves only as a correct-on-return guard for overflow/truncation (or implementation bugs). On fixed-width hardware, if a run ever leaves this regime, one may either take the conservative ClassicalMM fallback, or increase the staged-precision budget / word width and retry. Algorithm  2 formalizes a simple retry-with-more-staging wrapper, and Theorem  9.2 shows that it preserves the soft-quadratic bit/word complexity when the required precision $b(n)$ is polylogarithmic.
9.1 Retry with more staging (optional)
The ClassicalMM branch in Algorithm  1 is a correct-on-return backstop for implementations that leave the declared exact-execution regime (Appendix  B.5). In staged-precision fixed-word mode, a more efficient response is often to increase the staged mantissa budget and rerun the computation. The wrapper below does so without changing the kernel logic.
Attempt primitive.
Let KernelMM_Attempt(A,B;K)(A,B;K) denote Algorithm  1 executed in digit-band staged mode with $K$ words per scalar, except that the failure-policy fallback of Appendix  B.5 is replaced by returning a failure flag and aborting the attempt (i.e., no extracted value computed under a failed certificate is ever returned).
Input: $A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two; initial staging $K_{0}\geq 1$; maximum staging budget $K_{\max}$
Output: $C=AB$ exactly (or exact fallback if the staging limit is exhausted)
$K\leftarrow K_{0}$
while $K\leq K_{\max}$ do
 $(C,\mathrm{ok})\leftarrow\textsc{KernelMM\_Attempt}(A,B;K)$
 if $\mathrm{ok}$ then
    return $C$
 $K\leftarrow 2K$
return ClassicalMM(A,B)(A,B)
// correct-on-return backstop
[…]
Theorem 9.2 (Retry wrapper: correctness and complexity).
Fix a base word width $w$ and consider digit-band (slice) staging with $K$ words per scalar. Assume that for a full run of Algorithm  1 under $K$-word staging in the certified (no-failure) regime, the fixed-word cost satisfies $B(n,K)=\Theta(n^{2}\log n\cdot K)$ (each scalar primitive costs $\Theta(K)$ word operations).
[…]
Proof.
[…]
Correctness.
On the successful iteration, no node triggers the failure policy, so the attempted execution follows exactly the same computation path as Algorithm  1 and satisfies the hypotheses of Theorem  8.5, so it returns $C=AB$ exactly. If instead the wrapper exhausts the staging limit ($K>K_{\max}$), it returns ClassicalMM(A,B)(A,B), which is exact by Algorithm  3.
Cost.
For each failed attempt with budget $K_{i}$, the wrapper performs at most the work of a full staged run at the same $K_{i}$. Therefore the total cost is bounded by:
$\sum_{i=0}^{s}B(n,K_{i})~=~\Theta(n^{2}\log n)\sum_{i=0}^{s}K_{i}~=~\Theta(n^{2}\log n)\,(2K_{s}-K_{0})~=~O(n^{2}\log n\cdot K_{s}).$ (74)
Because $K_{s}<2K^{\star}$, this is $O(n^{2}\log n\cdot K^{\star})=O(B(n,K^{\star}))$, and likewise $O(B(n,2K^{\star}))$. The bound on the number of attempts follows from $K_{s}\geq K^{\star}$ and $K_{s}=2^{s}K_{0}$. ∎
@@ 9.2 Arithmetic recurrence
Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two and that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback), then:
@@
We now consider a single call of KernelMM on size $n$ under the assumption that the certificate-and-upgrade rule of Appendix  B.5 succeeds at this node and recursively at all descendants (so no fallback occurs).
@@ 9.3 Bit/word complexity (fixed words, staged precision)
Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers. Assume additionally that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback). Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
@@ 9.4 Example: integer polynomial convolution as a GPR-admissible packing
There exists a sufficient base sequence $\beta(\sigma)$ (growing with the node-local lattice exponent $\sigma$) such that the recursive packed convolution algorithm recovers $f\ast g$ exactly in dyadic lattice semantics. The arithmetic recurrence is:
@@
At each node, the packed product $F\ast G$ admits the three-band form (Proposition 9.10). Dyadic normalization (Lemma 9.9) ensures the band coefficients are bounded in the normalized lattice $\mathcal{L}_{\mathrm{coeff}}(\sigma)$. By choosing $\beta$ to satisfy the sufficient-base condition (Theorem 7.1) for the current node-local lattice exponent $\sigma$, the master gap inequality holds, and the scaled sieve $\mathcal{S}_{\beta,\sigma}$ recovers the middle band $T$ exactly. The algorithm performs $4$ subcalls (implicitly or explicitly to recover all components of the convolution from the packed forms), maintaining exactness throughout via the staged precision model.
@@ 10 Prior related work and what is new here
What is new in this note is not the existence of coefficient/band extraction, but the way recursive invariant amplification (centering plus dyadic $\ell_{2}$ normalization at every recursion node) makes coefficient extraction compatible with a model-honest intermediate-budget analysis: it gives a node-local dyadic base policy $\beta=\beta(\sigma)$ (Appendix B.5), eliminates bit-growth of mantissas, and improves feasibility/robustness by bounding intermediates at every depth rather than only at the root.
@@ 10.1 APA and border-rank methods in fast matrix multiplication
In fast matrix multiplication theory, approximate (APA) bilinear algorithms introduce a formal parameter (often written $\lambda$ or $\varepsilon$) to separate contributions across different powers of that parameter and then recover a desired coefficient via an exactification Step (interpolation/limit argument). This is similar to the work of Bini et al., on approximate MM [18] and in Bini’s later discussion relating exact and approximate bilinear algorithms [19]. A closely related tensor-language view is that such constructions upper-bound the border rank of the MM tensor, which is sufficient to derive asymptotic exponents (see [20] and Pan’s trilinear/field-extension methods [21]).
@@ 10.4 What changes the game: recursive invariant amplification
Part II Float-GPR: baseline comparators and commensurate accuracy
@@ 14 Float-GPR complexity
Part III Case studies and templates
@@ 17.3 How to create new GPR reductions
Part IV Generality of GPR: admissible packings, design patterns, and extensions
@@ 21.4 B) Beyond rounding gaps: an extractor abstraction
Proof.
Choose global packing parameters (e.g., bases $\beta$ or a fixed set of moduli) as in the hypothesis so that, for every depth $\ell$ and every packed value $Z$ produced at depth $\ell$, the uniform master bound:
$\|T_{\ell}\|_{N}\leq CS_{\ell}\leq CS_{\max}~\text{(or}~\|T_{\ell,E_{0}}\|_{N}\leq CS_{\ell}\leq CS_{\max}\text{)}$ (246)
implies $\mathsf{Cond}(Z)$. (The second inequality uses the definition of $S_{\max}:=\max_{\ell}S_{\ell}$.)
Extraction is correct at every node.
Fix any depth $\ell$ and any packed value $Z$ arising at depth $\ell$. By the preceding choice of parameters, $\mathsf{Cond}(Z)$ holds. By the extractor interface (Definition 21.8):
$\mathrm{Extract}(Z)~=~\mathrm{TargetPart}(Z),$ (247)
i.e., the extractor returns exactly the intended target coefficient block (degree band or multi-degree band).
Correctness propagates through the recursion.
We argue by structural induction over the recursion tree of the GPR reduction.
At a leaf node, the reduction evaluates the kernel on the (unpacked) leaf inputs directly and returns the result in the native domain; this equals the corresponding classical computation by definition.
For an internal node $u$ at depth $\ell$, let $(x_{1},\dots,x_{k})$ denote the node-local inputs in the native domain. The reduction forms packed embeddings:
$R_{i}(\boldsymbol{\beta})~=~\sum_{p\in D_{i}}\boldsymbol{\beta}^{p}R_{i}^{[p]}~~~(i=1,\dots,k)$ (248)
as in Definition 21.5, and computes the packed kernel value
$Z~:=~\mathcal{B}\left(R_{1}(\boldsymbol{\beta}),\dots,R_{k}(\boldsymbol{\beta})\right).$ (249)
By $k$-linearity of $\mathcal{B}$ (Definition 21.4), expanding gives:
$Z~=~\sum_{(p_{1},\dots,p_{k})\in D_{1}\times\cdots\times D_{k}}\boldsymbol{\beta}^{p_{1}+\cdots+p_{k}}~\mathcal{B}\left(R_{1}^{[p_{1}]},\dots,R_{k}^{[p_{k}]}\right).$ (250)
By construction of the packing scheme, the coefficient block(s) indexed by the target band (the degree or multi-degree window specified in the reduction) is exactly the value that the classical, unpacked computation at node $u$ is meant to return; this coefficient block is $\mathrm{TargetPart}(Z)$ by definition. By the first part of the proof, extraction at node $u$ is correct, so the reduction returns $\mathrm{Extract}(Z)=\mathrm{TargetPart}(Z)$, i.e., the correct classical node-local output. This completes the induction. Therefore the output returned at the root of the recursion coincides with the intended classical output.
Asymptotic complexity is preserved.
At a node operating on blocks of side length $m$, packing, extraction, and the associated bookkeeping (centering, normalization, digit/CRT/bucket operations) act entrywise and perform a number of primitive operations bounded by a constant $\kappa$ that depends only on the chosen global parameters (i.e., only on $C$ and $S_{\max}$ in the hypothesis), and not on $m$ or on the depth. Thus the nonrecursive work at a node is $O(\kappa\,m^{2})$, and the recursion has the same form as the baseline blocked recursion up to the multiplicative factor $\kappa$. For instance, under $2\times 2$ block recursion for matrix multiplication,
$T(m)~=~4T(m/2)+O(\kappa\,m^{2}),$ (251)
so $T(n)=O(\kappa\,n^{2}\log n)$. The same reasoning applies to bit/format complexity: the dependence on the packing/extractor parameters is absorbed into $\kappa$, and the asymptotic dependence on problem size follows the same recurrence. ∎
@@ B.4 Exact sieve extractor $\mathcal{S}_{\beta}(\cdot)$
Fix an integer $b\geq 1$ and let $\beta:=2^{b}$. Let $\sigma\geq 0$ be the node-local coefficient-lattice exponent at the current recursion node (Definition 3.2; Eq. 266). For $z\in\mathcal{L}_{D}$, define:
$\mathcal{S}_{\beta,\sigma}(z)$ $:=$ $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil$ (259)
$=$ $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}(z/\beta)\right\rceil.$ (260)
Here $z/\beta$ is implemented by exponent update $e\leftarrow e-b$ (exact in $\mathcal{L}_{D}$), and the rounding $\frac{1}{M}\left\lfloor M(\cdot)\right\rceil$ is implemented by rounding the mantissa to nearest integer at the combined exponent $e_{\mathrm{total}}=e+\sigma$.
Lemma B.7 (Implementation matches Part I sieve (with scaling)).
If $\mathcal{S}_{\beta,\sigma}$ is implemented as in Definition B.6, then it matches Definition 5.1 exactly, i.e.:
$\mathcal{S}_{\beta,\sigma}(z)~=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\,\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}(z/\beta)\right\rceil.$ (261)
In particular, Lemmas 5.25.3 apply verbatim to the implemented extractor.
@@
All steps are exact in $\mathcal{L}_{D}$ (i.e., dividing by $\beta=2^{b}$ is an exponent-offset update) and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma. Therefore the implemented $\mathcal{S}_{\beta}$ satisfies the mathematical sieve definition. ∎
B.5 Certified base choice and failure policy
Context and goal.
Here we give a fully deterministic, self-checking rule that makes an implementation of KernelMM (Algorithm 1) correct-on-return: before applying the sieve/extractor at a recursion node, the implementation certifies that the master-gap condition holds for the actual node-local lattice exponent $\sigma$ and the chosen packing base $\beta$. These checks are constant-cost (or reuse metadata already computed during normalization) and do not affect the near-quadratic scaling.
Fixed-slack sufficient-base condition (integral mode).
For the integral (exact dyadic) kernel we fix (once and for all) the slack constant:
$\delta_{0}~:=~2^{-3}=1/8,$ (262)
and we do not treat $\delta$ as a tuning parameter. At every recursion node we require the sufficient-base inequality from Theorem  7.1 with $\delta=\delta_{0}$:
$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta_{0},~~\beta=2^{b},~b\in\mathbb{Z}_{\geq 1}.$ (263)
When Eq. 263 holds and the normalization bounds $\|A^{\prime}_{i,:}\|_{2}\leq 1$, $\|B^{\prime}_{:,j}\|_{2}\leq 1$ are satisfied, extraction is exact because the master gap in Lemma  5.3 is open by at least $\delta_{0}$.
Tight node-local coefficient lattice exponent $\sigma$ (no over-estimation).
Represent each stored scalar as $x=(w/D)2^{e}$ (Appendix B.1) and write $\exp(x):=e$ for its exponent offset. After Step (2) of Algorithm 1 has produced normalized operands $A^{\prime},B^{\prime}$, define for each inner index $r\in\{1,\dots,m\}$ (where $m$ is the node dimension):
$a_{r}$ $:=$ $\min_{i}\exp(A^{\prime}_{ir}),$ (264)
$b_{r}$ $:=$ $\min_{j}\exp(B^{\prime}_{rj}),$ (265)
and let:
$g_{\min}$ $:=$ $\min_{r:a_{r}<\infty,~b_{r}<\infty}(a_{r}+b_{r}),$ (266)
$\sigma$ $:=$ $\max\{0,~-g_{\min}\}.$ (267)
with the convention that if the set $\{r:a_{r}<\infty,~b_{r}<\infty\}$ is empty (i.e., every dot-product term is zero) then $\sigma:=0$. This ensures that every dot-product term $A^{\prime}_{ir}B^{\prime}_{rj}$ lies in the scaled coefficient lattice:
$\mathcal{L}_{\mathrm{coeff}}(\sigma)~=~D^{-2}2^{-\sigma}\mathbb{Z}$ (268)
according to Definition 3.2.
Deterministic dyadic base selection with an exact integer certificate.
Let $A:=D^{2}2^{\sigma}\in\mathbb{Z}_{\geq 1}$ and $t:=\lfloor\log_{2}A\rfloor$. Because $A$ is just a power-of-two shift of the fixed integer $D^{2}$:
$t~=~\sigma+\left\lfloor\log_{2}(D^{2})\right\rfloor.$ (269)
Begin with the candidate exponent:
$b_{0}~:=~t+3~=~\sigma+\left\lfloor\log_{2}(D^{2})\right\rfloor+3,~~~\beta_{0}:=2^{b_{0}}.$ (270)
We now certify Eq. 263 for a dyadic base $\beta=2^{b}$ with $\delta_{0}=1/8$. After clearing denominators, the inequality Eq. 263 is equivalent to the integer inequality:
$8A(2^{b+1}+1)~\leq~3\cdot 2^{2b},$ (271)
where Eq. 271 can be evaluated exactly using integer arithmetic (shifts and adds).
One-step upgrade rule (guaranteed termination).
If Eq. 271 holds at $b=b_{0}$, set $b:=b_{0}$ and $\beta:=\beta_{0}$. Otherwise, set:
$b~:=~b_{0}+1,~~~\beta~:=~2^{b}.$ (272)
Now Eq. 263 holds automatically for this upgraded $\beta$, and because $2^{t}\leq A<2^{t+1}$ and $b=t+4$:
$A\left(\frac{2}{2^{b}}+\frac{1}{2^{2b}}\right)$ $\leq$ $2^{t+1}\left(\frac{2}{2^{t+4}}+\frac{1}{2^{2t+8}}\right)$ (273)
$=$ $\frac{1}{4}+2^{-(t+7)}$ (274)
$\leq$ $\frac{1}{4}+\frac{1}{128}$ (275)
$<$ $\frac{3}{8}~=~\frac{1}{2}-\delta_{0}.$ (276)
Note that the upgrade rule requires at most one doubling of $\beta$ and introduces no loops. In particular, in exact integer arithmetic the base certificate Eq. 271 must pass after at most one upgrade, i.e., any observed failure after the upgrade indicates overflow/truncation (or a bug) and triggers the failure policy below.
Failure policy (correct-on-return).
If the node certificate fails (either the normalization bounds are violated, or Eq. 271 fails even after the one-step upgrade), then the implementation must not apply the sieve at that node. Instead it must treat this as leaving the declared exact-execution regime (e.g., due to overflow, truncation below lattice resolution, or a bug) and must fall back to a conventional exact multiplication on the original subproblem, i.e., call ClassicalMM(A,B)(A,B) (Algorithm  3) and return its result. (An optimized implementation may instead increase the staged-precision budget / word width and restart the node; see Algorithm  2 and Theorem  9.2.) With this policy, the algorithm never returns an incorrect extracted value, i.e., failure affects only performance, not correctness.
Explicit fallback routine (ClassicalMM).
The fallback routine is the conventional triple-loop product evaluated in exact dyadic lattice semantics. A reference implementation may use a wider accumulator (or arbitrary-precision integers for the mantissas) for the inner sum, because this path is only a correctness backstop and is outside the near-quadratic performance model.
Input: $A,B\in\mathcal{L}_{D}^{m\times m}$
Output: $C=AB$ exactly
for $i\leftarrow 1$ to $m$ do
 for $j\leftarrow 1$ to $m$ do
    $s\leftarrow 0$
    for $r\leftarrow 1$ to $m$ do
       $s\leftarrow s+A_{ir}B_{rj}$
    $C_{ij}\leftarrow s$
return $C$
Algorithm 3 ClassicalMM(A,B)(A,B) (exact dyadic fallback)
Mandatory call-site (Algorithm 1).
In Algorithm 1 (KernelMM), compute $\sigma$ by Eq. 266 and set $\beta$ by the candidate Eq. 270 followed by the certificate-and-upgrade rule Eq. 271–Eq. 272.